OpenStudy (alexwee123):

Using polar equations, derive the formulas for the focal width of an ellipse and the focal width of a hyperbola.Begin by defining focal width for these conics in a manner analogous to the definition of the focal width of a parabola.

OpenStudy (dumbcow):

i don't get it, why do you need polar equations? the formula for the focal width is: ellipe c = sqrt(a^2 -b^2) hyperbola c = sqrt(a^2 +b^2)

OpenStudy (alexwee123):

i don't why too:/

OpenStudy (alexwee123):

but can you use polar to solve it please?

OpenStudy (experimentx):

I think put, \( \theta\) = 0 or \(\theta = \pi\) will give you r's, multiply it by 2\( e \) , that COULD be your focal width.

OpenStudy (dumbcow):

polar form for ellipse \[r = \frac{ab}{\sqrt{a^{2}\sin^{2} \theta + b^{2}\cos^{2} \theta}}\]

OpenStudy (alexwee123):

um

OpenStudy (alexwee123):

what about the directed angle?

OpenStudy (dumbcow):

not sure what you mean

OpenStudy (alexwee123):

theta

OpenStudy (alexwee123):

you told me how to get r but how would i get theta?

OpenStudy (dumbcow):

? polar equations are usually a function of theta ....r = f(theta) anyway from there i don't know how to derive the focal width c sorry

OpenStudy (alexwee123):

well thanks for all your help dumbcow andexperimentX

OpenStudy (experimentx):

 If the ellipse is standard ellipse centered on origin, then obviously, theta will be zero or pi

OpenStudy (dumbcow):

when theta = 0 or pi r = a

OpenStudy (experimentx):

Ah .. yes, 2ea <--- that will be our focal width. <--- should have thought earlier.